OpenStudy (anonymous):

help please 15x15 7/8

OpenStudy (anonymous):

its a fraction whole number x a fraction 15 x 15 7/8

OpenStudy (anonymous):

how did you do the problem

OpenStudy (anonymous):

yes i don,t know how to do this that why i am asking

OpenStudy (anonymous):

Is that what the problem is asking? 15 * 15 7/8 = 15 * (15*8 + 7)/8 = 15 * 127 / 8 = 1905/8

OpenStudy (pokemon23):

that's the question GT

OpenStudy (anonymous):

When you are asked such problems with a mixed fraction involved - like 15 7/8 - always convert that mixed fraction into an improper fraction first - like 127/8. Then, it is all basic arithmetic.

OpenStudy (anonymous):

it a word problem i write it down merchandise is stack 15 bxs high each bx is 15 7/8 how tall is the stack of bxs

OpenStudy (anonymous):

on the whole number how do you turn it into a fraction

OpenStudy (anonymous):

15 7/8 = 15 + 7/8 = (15*8 + 7)/8

OpenStudy (anonymous):

ok thank you